# Higher Lipschitz invariants

### v1

@@ Abstract.
We introduce new bi-Lipschitz invariants for functions of two complex variables, building on our earlier studies and on the foundational paper by Henry and Parusiński.
@@ 1. Introduction and preliminaries
Based on the clustering gradient canyons, we have shown in [PT] that the multi-level “identity card” of a 2-variable holomorphic function germ $f:({\mathbb{C}}^{2},0)\to({\mathbb{C}},0)$ is a discrete bi-Lipschitz-invariant. We have pointed out in [MPT] how the modular bi-Lipschitz invariant found by Henry and Parusiński [HP] can be embedded into this picture as the first of the coefficients in the expansion of $f$ along polars.
In this paper we find new bi-Lipschitz invariants of “HP-type” within the identity card of the clusters, toward a complete list of bi-Lipschitz invariants enabling us to decide the bi-Lipschitz type of a given two-variables holomorphic function germ. This contributes to a long standing research project, as one can move away from a merely homeomorphism equivalence (too weak) but not too close to a diffeomorphism equivalence (too rigid).
Our main results are Theorems 3.4, 3.5 and 3.6. Our Theorem 3.5 unveils a second level bi-Lipschitz invariant, based on two gradient canyons and their contact order. By Theorem 3.6 we show how to build, in a similar way, a third level bi-Lipschitz invariant of HP-type in case one disposes of 3 or more disjoint gradient canyons, and their contact orders satisfy certain conditions. As expected, this procedure may continue recursively toward still higher level, finitely many invariants of HP-type, cf Theorem 3.6.
Through Theorem 3.4 we infer a new and different spirit in the research of invariants. Let $f=g\circ\varphi$ for some bi-Lipschitz homeomorphism. Then Theorem 3.4 shows that, surprisingly, one may determine the development of $\varphi$ to a certain extent. The power of this result is illustrated by our second example in §4.
Our two examples have a key role in understanding two types of applications of Theorem 3.5. Example 4.1 shows how this theorem decides that any two generic function germs from a certain family are not bi-Lipschitz equivalent. In contrast, Theorem 3.5 is no longer conclusive in Example 4.2, where two generic function germs in a certain family turn out not to be bi-Lipschitz equivalent. To reach this conclusion we need to employ Theorem 3.4, which yields the existence of new bi-Lipschitz moduli at this level.
@@ 3.4. Higher invariants: second level
With these notations, we have:
Theorem 3.5.
@@
(20) $a_{1}y^{h}+\cdots+a_{1}\tilde{a}_{1}y^{H}+\cdots+O(y^{h+d_{1}-1})$ $=b_{1}Y_{1}^{h}+\cdots+\overline{b}_{1}y^{H}+\cdots+O(Y_{1}^{h+d_{1}-1}),$
$a_{2}y^{h}+\cdots+a_{2}\tilde{a}_{2}y^{H}+\cdots+O(y^{h+d_{2}-1})$ $=b_{2}Y_{2}^{h}+\cdots+\overline{b}_{2}y^{H}+\cdots+O(Y_{2}^{h+d_{2}-1}).$
@@
(22) $y^{h}+\cdots+\tilde{a}_{1}y^{H}+\cdots+O(y^{h+\delta-1})$ $=\frac{b_{1}}{a_{1}}Y_{1}^{h}+\cdots+\frac{\overline{b}_{1}}{a_{1}}y^{H}+\cdots+O(Y_{1}^{h+\delta-1}),$
$y^{h}+\cdots+\tilde{a}_{2}y^{H}+\cdots+O(y^{h+\delta-1})$ $=\frac{b_{2}}{a_{2}}Y_{1}^{h}+\cdots+\frac{\overline{b}_{2}}{a_{2}}y^{H}+\cdots+O(Y_{1}^{h+\delta-1}).$
@@
3.5. Higher level invariants: third level
@@ 4. Examples
Computing the invariants of Theorem 3.5 in the family of functions $F_{t}$ of Example 4.1 displayed below shows that, for small enough generic parameters $t$ and $s$, $F_{t}$ and $F_{s}$ are not bi-Lipschitz equivalent. In contrast, the invariants of Theorem 3.5 turn out to be not sufficient to settle the same question for the family $G_{t}$ in Example 4.2. It is by the key application of Theorem 3.4 that we can decide that $G_{t}$ and $G_{s}$ are generically not bi-Lipschitz equivalent.
@@
We will follow the proof of Theorem 3.5 in order to contradict the assumption that there exists a bi-Lipschitz homeomorphism $\varphi$ such that $F_{t}=F_{s}\circ\varphi$ for some parameters $s,t\in{\mathbb{C}}$ close enough to 00.
@@
where $d:=d_{\gamma_{1}}=d_{\gamma_{2}}=$, and where $Y_{1},Y_{2}$ are local variables, and $\gamma^{\prime}_{1}$, $\gamma^{\prime}_{2}$ are corresponding polar arcs of $F_{s}$. Since we have two options for the correspondence by $\varphi$ between the pair of canyons of $F_{s}$ and of $F_{t}$, the computation of the invariants of Theorem 3.5 falls into two cases.
@@
By Theorem 3.5, we must have $\frac{\tilde{b}_{1}-\tilde{b}_{2}}{\tilde{a}_{1}-\tilde{a}_{2}}=c^{h-H}$, which in our case becomes $\frac{s^{3}}{t^{3}}=c^{-3}$, while we have seen before that $c^{-12}=1$. The two equalities do not have a common solution $c$ This contradiction finishes our proof that, in Case 1, $F_{t}$ is not bi-Lipschitz equivalent to $F_{s}$ for a generic choice of values $t$ and $s$ of the parameter in our family.
@@
Pursuing the algorithm, we find (cf Theorem 3.5) that the effect of the bi-Lipschitz map on the pair $\bigl{(}H,(\tilde{a}_{1}-\tilde{a}_{2})\bigr{)}$ is the identity on $H$, and the multiplication of $(\tilde{a}_{1}-\tilde{a}_{2})$ by $c^{h-H}=c^{-\frac{3}{2}}=1$. Note that (32) tells that $c=1$. Up to this point, we do not get any contradiction, thus no proof or disproof of our assumption.

### v2

# On higher Lipschitz invariants

@@ Abstract.
We find new bi-Lipschitz invariants for functions of two complex variables.
@@ 1. Introduction and preliminaries
Based on the clustering gradient canyons, we have shown in [PT] that the multi-level “identity card” of a 2-variable holomorphic function germ $f:({\mathbb{C}}^{2},0)\to({\mathbb{C}},0)$ is a discrete bi-Lipschitz invariant. We have pointed out in [MPT] how the modular bi-Lipschitz invariant found by Henry and Parusiński [HP] can be embedded into this picture as the first of the coefficients in the expansion of $f$ along polars.
In this paper we infer a new and different spirit in the research of invariants. Let $f=g\circ\varphi$ for $f,g$ holomoprhic function germs of 2 variables, and some bi-Lipschitz homeomorphism $\varphi$. Our main result, Theorem 3.4, shows that, surprisingly, one may determine a development of $\varphi$ to a certain extent. The power of this result is illustrated by our second example in §4.
As a byproduct of our refined study [PT], we find here new bi-Lipschitz invariants within the identity card of the clusters, toward a complete list of bi-Lipschitz invariants enabling us to decide the bi-Lipschitz type of a given two-variables holomorphic function germ. This contributes to a long standing research project, as one can move away from a merely homeomorphism equivalence (too weak) but not too close to a diffeomorphism equivalence (too rigid). Our Corollary 3.5 unveils a second level bi-Lipschitz invariant of Henry-Parusiński type, based on two gradient canyons and their contact order. By Corollary 3.6 we show how to build, in a similar way, a third level bi-Lipschitz invariant of HP-type in case one disposes of 3 or more disjoint gradient canyons, and their contact orders satisfy certain conditions. As expected, this procedure may continue recursively toward still higher level, producing finitely many invariants of HP-type, cf Corollary 3.6.
Our two examples have a key role in understanding two types of applications of Corollary 3.5. Example 4.1 shows how this result decides that any two generic function germs from a certain family are not bi-Lipschitz equivalent. In contrast, Corollary 3.5 is no longer conclusive in Example 4.2, where two generic function germs in a certain family turn out not to be bi-Lipschitz equivalent. To reach this conclusion we need to employ our key Theorem 3.4, which yields the existence of new bi-Lipschitz moduli at this level.
Added in proof. After uploading our manuscript on arXiv, N. Nguyen pointed out to us his newly posted arxiv preprint [Ng] in which he observes a version of our Corollary 3.5 by deriving it directly from our constructions in [PT].
@@ 3.4. Higher invariants: second level
With these notations, we have [^2] After the publication of this manuscript as an arXiv preprint, N. Nguyen pointed up to us his manuscript [Ng] containing a statement which seems equivalent to our Theorem 3.5.:
Corollary 3.5.
@@
(20) $a_{1}y^{h}+\cdots+a_{1}\tilde{a}_{1}y^{H}+\cdots+O(y^{h+d_{1}-1})$ $=b_{1}Y_{1}^{h}+\cdots+\overline{b}_{1}Y_{1}^{H}+\cdots+O(Y_{1}^{h+d_{1}-1}),$
$a_{2}y^{h}+\cdots+a_{2}\tilde{a}_{2}y^{H}+\cdots+O(y^{h+d_{2}-1})$ $=b_{2}Y_{2}^{h}+\cdots+\overline{b}_{2}Y_{2}^{H}+\cdots+O(Y_{2}^{h+d_{2}-1}).$
@@
(22) $y^{h}+\cdots+\tilde{a}_{1}y^{H}+\cdots+O(y^{h+\delta-1})$ $=\frac{b_{1}}{a_{1}}Y_{1}^{h}+\cdots+\frac{\overline{b}_{1}}{a_{1}}Y_{1}^{H}+\cdots+O(Y_{1}^{h+\delta-1}),$
$y^{h}+\cdots+\tilde{a}_{2}y^{H}+\cdots+O(y^{h+\delta-1})$ $=\frac{b_{2}}{a_{2}}Y_{1}^{h}+\cdots+\frac{\overline{b}_{2}}{a_{2}}Y_{1}^{H}+\cdots+O(Y_{1}^{h+\delta-1}).$
@@
3.5. Higher invariants: third level
@@ 4. Examples
Computing the invariants of Corollary 3.5 in the family of functions $F_{t}$ of Example 4.1 displayed below shows that, for small enough generic parameters $t$ and $s$, $F_{t}$ and $F_{s}$ are not bi-Lipschitz equivalent. In contrast, the invariants of Corollary 3.5 turn out to be not sufficient to settle the same question for the family $G_{t}$ in Example 4.2. It is by the key application of Theorem 3.4 that we can decide that $G_{t}$ and $G_{s}$ are generically not bi-Lipschitz equivalent.
@@
We will follow the proof of Corollary 3.5 in order to contradict the assumption that there exists a bi-Lipschitz homeomorphism $\varphi$ such that $F_{t}=F_{s}\circ\varphi$ for some parameters $s,t\in{\mathbb{C}}$ close enough to 00.
@@
where $d:=d_{\gamma_{1}}=d_{\gamma_{2}}$, and where $Y_{1},Y_{2}$ are local variables, and $\gamma^{\prime}_{1}$, $\gamma^{\prime}_{2}$ are corresponding polar arcs of $F_{s}$. Since we have two options for the correspondence by $\varphi$ between the pair of canyons of $F_{s}$ and of $F_{t}$, the computation of the invariants of Corollary 3.5 falls into two cases.
@@
By Corollary 3.5, we must have $\frac{\tilde{b}_{1}-\tilde{b}_{2}}{\tilde{a}_{1}-\tilde{a}_{2}}=c^{h-H}$, which in our case becomes $\frac{s^{3}}{t^{3}}=c^{-3}$, while we have seen before that $c^{-12}=1$. The two equalities do not have a common solution $c$ This contradiction finishes our proof that, in Case 1, $F_{t}$ is not bi-Lipschitz equivalent to $F_{s}$ for a generic choice of values $t$ and $s$ of the parameter in our family.
@@
Pursuing the algorithm, we find (cf Corollary 3.5) that the effect of the bi-Lipschitz map on the pair $\bigl{(}H,(\tilde{a}_{1}-\tilde{a}_{2})\bigr{)}$ is the identity on $H$, and the multiplication of $(\tilde{a}_{1}-\tilde{a}_{2})$ by $c^{h-H}=c^{-\frac{3}{2}}=1$. Note that (32) tells that $c=1$. Up to this point, we do not get any contradiction, thus no proof or disproof of our assumption.